Some eighth-grade students want to raise at least $300 for a field trip by selling popcorn and fruit bars.
The chart below shows the amount of profit they will make on each sale.
If they sell exactly 500 fruit bars, how many boxes of popcorn will they need to sell to make a total of $300?
popcorn=60 cents
fruitbars=30 cents

Answer by addingup(3677)

Amount to raise: 300
popcorn = 60 cts
fruitbars = 30 cts
.
they sold 500 fruit bars = 500 x 0.30 = 150
.
(300 - 150)/0.6 = 150/0.6 = 250 this is how many popcorn they need to sell